Could you show me the steps you take to solve:
The sum of two integers is thirty. Eight times the smaller integer is six more than five times the larger integer. Find the integers.

Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

The sum of two integers is thirty.
x + y = 30
------------------
Eight times the smaller integer is six more than five times the larger integer.
8x = 5y +6
----------------
Arrange for elimination:
8x + 8y = 240
8x - 5y = 6
---------------
Find the integers.
------------------------
Subtract the 2nd equation from the 1st to get:
13y = 234
y = 18 (larger integer)
----------
Substitute into x + y = 30 to solve for "x":
x + 18 = 30
x = 12 (smaller integer)
-------------------
Cheers,
Stan H.





Answer by nerdybill(7384)

The sum of two integers is thirty. Eight times the smaller integer is six more than five times the larger integer. Find the integers.
.
Assign variables to things you don't know.
Let x = larger integer
and y = smaller integer
.
Since we have two unknowns, we'll need to find two equations.
From: "The sum of two integers is thirty." we get equation 1:
x+y = 30
.
From: "Eight times the smaller integer is six more than five times the larger integer." we equation 2:
8y = 5x+6
.
To solve, let's solve equation 1 for 'x':
x+y = 30
x = 30-y
.
Substitute the above into equation 2 and solve for y:
8y = 5x+6
8y = 5(30-y)+6
8y = 150-5y+6
13y = 150+6
13y = 156
y = 156/13
y = 12 (value of smaller integer)
.
Substitute the above into equation 1 and solve for x:
x+y = 30
x+12 = 30
x = 30-12
x = 18 (value of larger integer)